The Memorandum of Agreement (MOA) that allows us to complete the construction and installation of the FEL Upgrade was signed this week by DOE and ONR. This MOA will allow $4.5M of FY2001 DOD funds to be released to Jefferson Lab for the FEL program. We thank Eli Zimet, John Albertine and Wayne Skinner for their efforts and continuing support of the ONR/DOE collaboration which makes the FEL program possible.
We are also pleased to announce that the FEL program won a $400k grant from DOD to study the effects of femtosecond, pulsed and cw laser irradiation on laser-materials interactions.
The project received an invitation from the Defense Science Board subcommittee on high energy lasers to present a status report on the Jefferson Lab FEL program on Jan. 25.
WBS 3 (Beam Physics):
Indecision on final lattice parameters, driven by the designer’s concern about risk management and his work arounds for the ongoing lack of adequate design tools, continued. Approximate specifications for various elements were given to WBS 9.
WBS 5 (SRF):
Jefferson Lab machine shop has completed additional tooling for cavity cell machining which will significantly increase cavity production rate. Helium Vessel vendor, PHPK, has detailed a modification to the helium vessel and will send samples of new weld next week. All other contracts are on schedule.
WBS 8 (Instrumentation):
The major effort is upgrading the entire Laser Safety System (LSS) to accept the use of an enclosed hutch. This change will make all labs respond/certify/logic the same. By inserting and turning a key in the local LSS box one can enable a (multiple) hutch. Additionally six more monitors were added to the control room dedicated to monitoring the user labs, this couples with "stop light" status lamps allow an instant view of all user lab activities. On the individual user lab doors stop lights were added to inform the user of the following conditions: RED - NO Access, YELLOW - Alignment Mode; Goggles Required, GREEN - Authorized users holding a valid smart card may enter.
The alignment mode was enhanced to give the operator an option of running 60Hz, 10 microsecond macropulses, or the standard 2 Hz, 250 microsecond macropulses. Both of these activities were considerable more difficult to realize that the words to describe them...
WBS 9 (Transport):
The task of designing the Injection and Extraction dipoles (GU & GV) was awarded to DULY Research. G. Biallas is going there early next week to kick off the design.
The task of designing the UV Modified GX, the Reverse Bend (GQ) and the 180 degree bend (GY) was awarded to Advanced Energy Systems. In the same trip as above, G. Biallas will kick off the design there.
Dave Douglas was able to issue more solid specifications for all the magnets of the engineering version of the lattice. We worked on making sure there is enough gap in the magnets to accept the thick vacuum chambers necessary to span the wide good field width required by the electron beam as well as have enough aperture to not vignette the expanding optical beam near the mirrors.
3 inch bore magnets
We evaluate a simpler method of construction for yokes that was suggested by a vender. When considering the entire
system with requirements for tooling balls for alignment, placement on the magnet measurement stand and final attachment to
the girders, we decided that the simple method was not so simple after all and would cost the same. We decided to stick with
our present design.
Trim Quad for the Arcs
Tom Hiatt, (an engineer recently freed up from SNS work to work half time) and Jeff Karn started forming the concept of a
real magnet from the specifications we received from Dave Douglas.
Measurement Probe: The 50 and 100 turn litz coils passed tests
for quality and continuity. The fiberglass plastic parts for the
probe body were cut by the vendor to rough shape and delivered back to us for annealing.
Sextupoles and Octupoles
With the advent of the firm specifications, Robin Wines was able to
start on forming the concept of both magnets using the literature search
a study she performed earlier. We will forgo making a combined Corrector
/ Sextupole - Octupole in favor of the more conservative and easily analyzed
Sextupole - Corrector combination.
Path Length Correctors
Dave Douglas was able to give specifications for the correctors that were just beyond the ends of the 180° bends in the IR Demo. We envision combining them into the ends of the 180° bends on the Upgrade, getting the return yoke for free, placing the field in a more ideal position and freeing up space on the beam tube between magnets for the Octupoles.
We ordered flanges for the shielded bellows as well as the unshielded bellows for the vacuum chamber.
The layout of the engineering version of the lattice progressed with the addition of preliminary dipole icons and the redefinition of the configuration of the change in the GX dipoles to switch out the UV electron beam.
WBS 10 (Wiggler):
Assembly of the final configuration of the wigglers started: making up the water manifolds, their mounting brackets and mounting the new pole tip clamps to achieve the required gap.
Tom Hiatt visited the manufacturer of the dispersion section coils just as they were about to start winding coils. The steel for the magnet was rough machined and sent for annealing.
PECO is just starting the final machining on the magnet poles and has completed the magnet base. We expect the magnet to be complete by the end of the month. All remaining drawings for the wiggler viewer optical transport and the vacuum chamber are in check. The drawings for the dispersion section support are signed off. The order for the wiggler beryllium viewers is out for bid. Good progress is being made on the assembly of the water manifold of the optical klystron.
WBS 11 (Optics):
We received the final report for our contract at AES on the deformable high reflecting mirror over the holiday, and have been reviewing it. In particular, we have to perform a careful check on all drawings before they can be approved for release. This has been the major activity that occurred this week. In the meantime, preliminary copies of the drawings have been sent out for bid. We have received some of our optical quality standards, as well as more optical components.
No operations this week. We will commence some pre-ops testing next
week when the LCW system has been repaired.